OpenStudy (anonymous):

Find all solutions (x, y) of the given system, where x and y are real numbers. (If there is no real solution, enter NO REAL SOLUTION.) y= x + 7 y= 49 − x^2 I found the smaller x-value which is (-7,0) I need help with finding the larger x-value?

OpenStudy (anonymous):

u need to equate both as following x+7=49-x^2 or (x+7)-(x^2-7^2)=0 or(x+7)-(x-7)(x+7)=0 or(x+7(1-x+7)=0 or (x+7)(8-x)=0 hence x=-7 or x=8 if x=-7 y=0 if x=8 y=7+8=15 (-7,0) and (8,15) are the reqd solutions